# Comparative Performance Analysis of Quantum Machine Learning Architectures for Credit Card Fraud Detection

### v1

@@ VI Conclusion
This study provides an insightful comparison of the performance of VQC, SQNN, and EQNN models in the context of financial fraud detection, highlighting their unique capabilities and limitations. By systematically exploring the impact of feature maps and ansatz configurations across two diverse datasets, the analysis underscores the critical factors influencing the effectiveness of QML models. The results demonstrate that VQC and SQNN models are highly competitive in their ability to detect fraud, with VQC achieving the highest performance. In contrast, EQNN showed limited capacity to handle the complexity of the datasets, pointing to potential areas for architectural enhancement. The findings emphasize the significance of tailoring model configurations, including selecting feature maps and ansatzes, to align with specific data characteristics and performance requirements.
This work contributes to a deeper understanding of the interplay between quantum model design and application context, providing a foundation for optimizing QML models in financial fraud detection. Future research can build on these findings by exploring advanced quantum architectures and alternative configurations to further enhance the utility of QML in tackling fraud and other high-stakes classification problems.
@@ Acknowledgment
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008.

### v3

@@ VI Conclusion
This study provides an insightful comparison of the performance of VQC, SQNN, and EQNN models in the context of financial fraud detection, highlighting their unique capabilities and limitations. By systematically exploring the impact of feature maps and ansatz configurations across two diverse datasets, the analysis highlights the critical factors that impact the effectiveness of QML models. The results demonstrate that VQC and SQNN models are highly competitive in their ability to detect fraud, with VQC achieving the highest performance. In contrast, EQNN shows limited capacity to handle the complexity of the datasets, pointing to potential areas for architectural enhancement. The findings emphasize the significance of tailoring model configurations, including selecting feature maps and ansatz, to align with specific data characteristics and performance requirements.
Beyond comparative evaluation, this work makes two complementary contributions. First, a one-way ANOVA test verifies the statistical significance of the effects of model structure, dataset, feature map, and ansatz on performance. This adds methodological rigor and confirms that the observed differences are not due to random variation. Second, the robustness of the best-performing model for each dataset is evaluated under five types of quantum noise. The results show that they maintain acceptable performance levels under noisy conditions, supporting their potential deployment on NISQ devices.
Future research can build on these findings by exploring more expressive quantum architectures, deeper circuits, and improved feature encodings. The performance limitations observed in some models may result from factors such as limited circuit depth, insufficient ansatz expressiveness, or mismatches between configuration and data complexity. Further investigation is needed to isolate these factors and guide the development of more effective QML solutions for complex classification tasks.
@@ Acknowledgment
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008, and the Center for Cyber Security (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104.